a science class has a ratio of 3 boys to 2 girls.what is the largest class that could be formed with less than 40 students?

Answer by ewatrrr(24785)

 
Hi,
a science class has a ratio of 3 boys to 2 girls.
B/G = 3/2 | Cross Multiplying
3G = 2B
G = (2/3)B
Question states***
what is the largest class that could be formed with less than 40 students?
B + (2/3)B = 40 for ex
B = (3/5)*40 18 and 12
B = (3/5)35 21 and 14 35 the largest class to maintain the ratio